OpenStudy (anonymous):

need help complete the ordered triple 4a+3b+c=-9 a-3b+2c=27 11a-2b+3c=28

OpenStudy (anonymous):

you add frist and second or add frist to thrid or second to thrid

OpenStudy (anonymous):

i think your suppose to multiply (1) by -2 then add to (2) which then becomes (4) then multiply (1) by -3 then add to (3) which becomes (5) then multiply (5) by-7 and then add to (4) then substitute in (5) then subtitute for (1) is this right?

OpenStudy (anonymous):

depen some time negative side only ,some time you time negatine side and number , some you don't have to time any thing just add. I will slove and get back to you when I ready

OpenStudy (anonymous):

ok please do

OpenStudy (anonymous):

add frist and second equation you just add , don't need time

OpenStudy (anonymous):

so that would be 5a+3c=18

OpenStudy (anonymous):

4a+3b+c=-9 a-3b+2c=27 add.................................... .................5a +ob +3c =18 you don't have to write ob or: 5a +3c =18 you don't have to write ob

OpenStudy (anonymous):

yes you are right

OpenStudy (anonymous):

now you can add; 3rd + 2nd or ; 1st = 3rd which one you want/

OpenStudy (anonymous):

let do 1st + 3rd ore easy

OpenStudy (anonymous):

my computer is fozen I have close and open internet again I lose every thing I solve

OpenStudy (anonymous):

Let do again

OpenStudy (anonymous):

ok 4a+3b+c=-9 a-3b+2c=27 11a-2b+3c=28

OpenStudy (anonymous):

Let do 1st+ 3rd First equation Time (-2) : 4a+3b+c=-9 ………………………-8a-6b-2c=18 Third equation Time (3) : 11a+2b+3c=28 ……………………… 33a +6b +9c = 84

OpenStudy (anonymous):

Let do 1st+ 3rd First equation Time (-2) : ......................4a+3b+c=-9 ………………-8a-6b-2c=18 Third equation Time (3) : ................... 11a+2b+3c=28 ……………… 33a +6b +9c = 84

OpenStudy (anonymous):

sorry 11a+2b+3c=18 (not 28) some time I tye wrong

OpenStudy (anonymous):

my mind not feel good,give me more time

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Ok. I done let me scan send can see better give me few more minute

OpenStudy (anonymous):

ok.you go this link http://www.flickr.com/photos/62007598 @N03/5645382314/

OpenStudy (anonymous):

I'm not good with tye , writing more easy for me